# Learning to Control Unknown Strongly Monotone Games

### v1

@@ 3.1 Performance Guarantees
We make the following assumption on our game to ensure that $\mathcal{N}_{opt}$ is non-empty, which is is just Slater’s condition [63] for convex optimization.
@@
The action set is of the form $\mathcal{X}=\{\boldsymbol{x}\mid h_{i}(\boldsymbol{x})\leq 0,i=1,\ldots,M\}$ for some $M$ where each $h_{i}(\boldsymbol{x})$ is a continuously differentiable convex function and there exists $\boldsymbol{x}\in\text{int}(\mathcal{X})$ (i.e., $h_{i}(\boldsymbol{x})<0$, $i=1,\ldots,M$) such that $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$.
We also take the following standard assumptions on our stepsize sequences [16].
@@
The stepsizes $\{\eta_{t}\}$ and $\{\epsilon_{t}\}$ are of the form $\eta_{t}=1/(t+T_{1})^{\eta}$ and $\epsilon_{t}=1/(t+T_{2})^{\epsilon}$ which satisfy:
@@
Under Assumptions 1-5, $\boldsymbol{\alpha}_{t}$ converges to the set $\mathcal{N}_{opt}$, $\boldsymbol{x}_{t}$ converges to $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$ and
@@
Therefore the best rate is obtained for $\eta=0.5$ and $\epsilon=0.75$:
$\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq O\left(t^{-1/4}\right).$
@@ 4 Convergence Analysis
In this section, we detail our proof strategy for Theorem 1 and 2. The proofs for the lemmas below and the main theorems are given in Appendix I and II respectively.
Throughout the proofs, we use the notation $F(\boldsymbol{x},\boldsymbol{\alpha})=F(\boldsymbol{x})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}$, which is the gradients vector of the game given the controlled input $\boldsymbol{\alpha}$.
Our first lemma shows that the desired set $\mathcal{N}_{opt}$ is non-empty. The proof for this lemma follows from the strongly monotone nature of the game and Slater’s conditions (Assumption 4).

### v2

@@ 3.1 Performance Guarantees
We make the following assumption on our game to ensure that $\mathcal{N}_{opt}$ is non-empty, which is just Slater’s condition [61] for convex optimization.
@@
The action set is of the form $\mathcal{X}=\{\boldsymbol{x}\mid h_{i}(\boldsymbol{x})\leq 0,i=1,\ldots,M\}$ for some $M$ where each $h_{i}(\boldsymbol{x})$ is continuously differentiable and convex and there exists $\boldsymbol{x}\in\text{int}(\mathcal{X})$ (i.e., $h_{i}(\boldsymbol{x})<0$, $i=1,\ldots,M$) such that $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$.
We also make the following standard assumptions on our step-size sequences [11].
@@
The stepsizes $\{\eta_{t}\}$ and $\{\epsilon_{t}\}$ are of the form $\eta_{t}=1/(t+T_{1})^{\eta}$ and $\epsilon_{t}=1/(t+T_{2})^{\epsilon}$ for some $T_{1},T_{2}>0$ which satisfy:
@@
Under Assumptions 1-5, $\boldsymbol{\alpha}_{t}$ converges to the set $\mathcal{N}_{opt}$, $\boldsymbol{x}_{t}$ converges to $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$, and
@@
Therefore the best rate is obtained for $\eta=0.5+\delta/3$ and $\epsilon=0.75+\delta$:
$\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq O\left(t^{-0.25+\delta}\right),$
where $\delta$ is arbitrarily small.
@@
A smaller value of $\epsilon$ gives a better rate. The optimal rate is then dictated by the assumptions that $\eta_{t}$ is square-summable ($\eta>0.5$) and $\lim_{t\uparrow\infty}\epsilon_{t}^{2}/\eta_{t}^{3}=0$ ($2\epsilon>3\eta$).
@@ 4 Convergence Analysis
In this section, we detail our proof strategy for Theorem 1 and 2. Throughout the proofs, we denote the gradient vector given the controlled input $\boldsymbol{\alpha}$ by $F(\boldsymbol{x},\boldsymbol{\alpha})=F(\boldsymbol{x})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}$.
Our first lemma shows that the desired set $\mathcal{N}_{opt}$ is non-empty. The proof follows from the strong monotonicity of the game and Slater’s conditions (Assumption 4).
@@
Proof.
Let $\mathcal{G}(\boldsymbol{\alpha})$ be our $N$-player game with utility functions $u_{n}(\boldsymbol{x},\boldsymbol{\beta})$ and action set $\mathcal{X}$ where $\boldsymbol{\beta}=\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}$. Consider a modified $N$-player game $\mathcal{G}^{\prime}$ with utility function $u^{\prime}_{n}(\boldsymbol{x})=r_{n}(\boldsymbol{x})$ for $n\in\{1,\ldots,N\}$ and action set $\{\boldsymbol{x}\mid\boldsymbol{x}\in\mathcal{X},\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}\}$. Under the assumption that there exists $\boldsymbol{x}$ such that $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$ and $h_{i}(\boldsymbol{x})<0$ for all $i=1,\ldots,M$, Theorem 1 along with equations (3.3) and (3.6) from [8] states that there exists a NE $\hat{\boldsymbol{x}}\in\mathcal{X}$ of $\mathcal{G}^{\prime}$, and $\boldsymbol{\mu}\in\mathbb{R}^{M},\boldsymbol{\lambda}\in\mathbb{R}^{K}$ such that
$F(\hat{\boldsymbol{x}})+\sum_{i=1}^{M}\nabla h(\hat{\boldsymbol{x}})\mu_{i}+\boldsymbol{A}^{\intercal}\boldsymbol{\lambda}=0$
and
$\boldsymbol{A}\hat{\boldsymbol{x}}=\boldsymbol{\ell^{*}}\;\textrm{and}\;h_{i}(\hat{\boldsymbol{x}})\leq 0,\forall i\in\{1,2,\ldots,M\}.$
Let $\boldsymbol{\alpha}=-\boldsymbol{\lambda}$. Then the above equations can be written as
$F(\hat{\boldsymbol{x}},\boldsymbol{\alpha})+\sum_{i=1}^{M}\nabla h(\hat{\boldsymbol{x}})\mu_{i}=0$
and
$h_{i}(\hat{\boldsymbol{x}})\leq 0,\forall i\in\{1,2,\ldots,M\},$
which implies that $\hat{\boldsymbol{x}}$ is a NE for the game $\mathcal{G}(\boldsymbol{\alpha})$, and hence $\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\hat{\boldsymbol{x}}$. We also know that $\boldsymbol{A}\hat{\boldsymbol{x}}=\boldsymbol{\ell^{*}}$. Hence there exists $\boldsymbol{\alpha}$ such that $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$, showing that $\mathcal{N}_{opt}$ is non-empty.
@@ 4.1 Non-expansive Mapping
Proof.
For a fixed $\boldsymbol{\alpha}$, $-F(\boldsymbol{x},\boldsymbol{\alpha})$ is strongly monotone on $\mathcal{X}$ and is Lipschitz continuous by Assumptions 1 and 2 respectively. Then [51, Theorem 2.3.3] states that for all $\boldsymbol{x}\in\mathcal{X}$, for some $L^{\prime}$
$\|\boldsymbol{x}-\boldsymbol{x^{*}}(\boldsymbol{\alpha})\|\leq L^{\prime}\|\boldsymbol{x}-\Pi_{\mathcal{X}}(\boldsymbol{x}+F(\boldsymbol{x},\boldsymbol{\alpha}))\|.$
Hence for $\boldsymbol{\alpha}_{2},\boldsymbol{\alpha}_{1}\in\mathbb{R}^{K}$, we have
$\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|$
$\leq L^{\prime}\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{1}))\|$
=$\stackrel{{\scriptstyle(a)}}{{=}}L^{\prime}\Big{\|}\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2}))$
$\;\;\;\;\;\;\;\;\;-\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{1}))\Big{\|}$
$\leq L^{\prime}\|F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2})-F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{1})\|$
$=L^{\prime}\|\boldsymbol{A}^{\intercal}(\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1})\|$
$\leq L_{0}\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|,$
where $L_{0}=L^{\prime}\|\boldsymbol{A}^{\intercal}\|$. Here equality (a) follows from [51, Proposition 1.5.8] which says
$\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})=\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2})).$
This completes the proof for part (a) of Lemma 4.
Now let $\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2}\in\mathbb{R}^{K}$ Since $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})$ is the NE corresponding to $\boldsymbol{\alpha}_{1}$, using [51, Proposition 1.4.2], we have for every $\boldsymbol{x}\in\mathcal{X}$ that $\left\langle\boldsymbol{x}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{1}\right)\right\rangle\leq 0$. For $\boldsymbol{x}=\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})$
$\left\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{1}\right)\right\rangle\leq 0.$
Similarly, we also have
$\left\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2}\right)\right\rangle\leq 0.$
By combining the above two inequalities we obtain
$\left\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2}\right)-F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{1}\right)\right\rangle\geq 0.$ (11)
Then
$-\mu\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2}$
$\stackrel{{\scriptstyle(a)}}{{\geq}}\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{1}\right)-F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{1}\right)\rangle$
$=\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2}\right)-F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{1}\right)\rangle$
$\;\;\;+\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{2}-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{1}\rangle$
$\stackrel{{\scriptstyle(b)}}{{\geq}}\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{2}-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{1}\rangle$
$=(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}))^{\intercal}\boldsymbol{A}^{\intercal}(\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1})$
$=(\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}))^{\intercal}(\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1})$
$=\left\langle\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\right\rangle.$
Here inequality (a) uses that $F(\boldsymbol{x},\boldsymbol{\alpha})$ is strongly monotone in $\boldsymbol{x}$ with parameter $\mu$ and (b) uses (11). ∎
@@
Proof.
For any $\boldsymbol{\alpha}_{2},\boldsymbol{\alpha}_{1}\in\mathbb{R}^{K}$,
$\|g(\boldsymbol{\alpha}_{2})-g(\boldsymbol{\alpha}_{1})\|^{2}$
$=\|\boldsymbol{\alpha}_{2}+\gamma\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{\alpha}_{1}-\gamma\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2}$
$=\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|^{2}+\gamma^{2}\left\|\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\right\|^{2}$
$\;\;\;+2\gamma\left\langle\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\right\rangle$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|^{2}+\gamma^{2}\|\boldsymbol{A}\|^{2}\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2}$
$\;\;\;-2\mu\gamma\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2}$
$\leq\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|^{2}-(2\mu\gamma-\gamma^{2}\|\boldsymbol{A}\|^{2})\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|^{2}.$
Here inequality (a) follows from part (b) of Lemma 4 and inequality (b) follows from the assumption on $\gamma$ that $2\mu\gamma>\gamma^{2}\|\boldsymbol{A}\|^{2}$. Hence $g(\cdot)$ is non-expansive. ∎